prettygirl15:

y= 2x^2 + 16x + 34 Vertex Value of a Opens

xxDeppressionxx:

Step by Step Solution More Icon Rearrange: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : y-(2*x^2+16*x+34)=0 STEP1: Equation at the end of step 1 y - ((2x2 + 16x) + 34) = 0 STEP2: Equation at the end of step 2 y - 2x2 - 16x - 34 = 0 STEP3: Solving a Single Variable Equation: 3.1 Solve y-2x2-16x-34 = 0 (Source)= https://www.tiger-algebra.com/drill/y=2x~2_16x_34/

xxDeppressionxx:

Picture from https://www.mathpapa.com/algebra-calculator.html

xxDeppressionxx:

xxDeppressionxx:

That’s if you need to solve that equation.

prettygirl15:

thx

xxDeppressionxx:

Np